OpenStudy (anonymous):

Choose the two equations you would use to solve the absolute value equation below. Then solve the two equations. |x – 7| = 12 A. x + 7 = –12 and x + 7 = 12; {–19, 5} B. x + 7 = –12 and x + 7 = 12; {–19, 19} C. x – 7 = 12 and x – 7 = –12; {19, –5} D. x – 7 = 12 and x – 7 = –12; {–19, 19}

jimthompson5910 (jim_thompson5910):

hint: if |x| = k, then x = k or x = -k where k is some positive number

jimthompson5910 (jim_thompson5910):

example if |x| = 3, then x = 3 or x = -3

OpenStudy (anonymous):

|a+b| = c Solved as a +b = c and a +b = -c

OpenStudy (anonymous):

Is it A?

jimthompson5910 (jim_thompson5910):

Similar Example (just with different numbers) if |x-19| = 4 then x-19 = 4 or x-19 = -4

jimthompson5910 (jim_thompson5910):

from there, you solve each equation for x

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Did you check your answers by plugging them back in the equation?

OpenStudy (anonymous):

Yes the answers A

jimthompson5910 (jim_thompson5910):

unfortunately that is incorrect

OpenStudy (anonymous):

How?

jimthompson5910 (jim_thompson5910):

Notice how |x – 7| = 12 has a minus between the x and 7

jimthompson5910 (jim_thompson5910):

but choice A has a plus between the x and 7

OpenStudy (anonymous):

So its C

OpenStudy (anonymous):

|x-7| = 12 to solve we would say x-7 = 12 and x -7 = -12 solve for x for both equations

jimthompson5910 (jim_thompson5910):

much better

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yay :D lol thanks